OpenStudy (anonymous):

what is the slope and y-intercept of the line represented by the equation below? 13x-8y=12 A.slope=13/8 y-intercept=-3/2 B.slope=-13/8 y-intercept=3/2 C.Slope=13/8 y-intercept=3/2 D.Slope=-13/8 y-intercept=-3/2

OpenStudy (anonymous):

please help me!

OpenStudy (luigi0210):

The y-inter is when x=0, so that one should be fairly easy to find.

OpenStudy (anonymous):

im not sure how to do it

OpenStudy (anonymous):

help me get the answer?

OpenStudy (luigi0210):

Do you see where the x is? Plug a 0 in there.

OpenStudy (anonymous):

k then what?

OpenStudy (luigi0210):

Solve for y.

OpenStudy (anonymous):

i dont get it is okay tho

OpenStudy (luigi0210):

Okay here: To solve for the y-inter we plug in for x: 13(0)-8y=12 0-8y=12 Now we divide by -8 y=-12/8 Simplify it down to get: y=-3/2 So we know the answer has to have a -3/2 eliminating half the choices

OpenStudy (anonymous):

okay thank you!

OpenStudy (luigi0210):

Do you know hot to find the slope?

OpenStudy (anonymous):

well no i think i do but i might be missing something

OpenStudy (anonymous):

would it be D?

OpenStudy (luigi0210):

Try it and show me what you get

OpenStudy (luigi0210):

D is wrong

OpenStudy (luigi0210):

13x-8y=12 -8y=-13x+12 y=13x/8+3/2

OpenStudy (anonymous):

is that for the slope?

OpenStudy (luigi0210):

Yes

OpenStudy (luigi0210):

So the answer by default is A

OpenStudy (luigi0210):

*y=13x/8-3/2

OpenStudy (anonymous):

ok thank you so much

OpenStudy (luigi0210):

Yup!